OpenStudy (anonymous):

2(h-8)-h=h-16 solve the equation? Ill give medal!

OpenStudy (***[isuru]***):

u cant solve it... it's the same equation u got on both side.. 2(h-8)-h=h-16 2(h-8) = 2h -16 2(h-8) = 2(h-8) C ??? Hope this will help ya !!!